On January 1, 2021, Maywood Hydraulics leased drilling equipment from Aqua Leasing for a three-year period ending December 31, 2024, at which time possession of the leased asset will revert back to Aqua. The equipment cost Aqua $416,676 and has an expected economic life of four years. Aqua expects the residual value at December 31, 2024, to be $54,000. Negotiations led to Maywood guaranteeing a $76,000 residual value.

Equal payments under the lease are $108,000 and are due on December 31 of each year with the first payment being made on December 31, 2021. Maywood is aware that Aqua used a 6% interest rate when calculating lease payments. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

Required:

1. & 2. Prepare the appropriate entries for Maywood on January 1, 2021 and December 31, 2021, related to the lease. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field. Round your final answers to the nearest whole dollar.)
Answer #1
This lease is consider as finance lease. Because of the lease period is equal to 75% (3 years lease / 4 expected life).
Year PV factor @6%
1 0.94340
2 0.89000
3 0.83962
Total PV annuity due (Total of 0 to 4) 2.67301
Present value of lease payments (108000*2.67301) $     288,685
Present value of probable residual payment due (76000-54000= 22000) (22000*0.83962) $       18,472
Total present value $     307,157
Lease amortization table
Date Beginning balance of lease payable Lease payment Interest expense (beginning balance of lease payable *11% ) Reduction in lease payable (lease payment - interest expense) Ending balance of lease payable (beginning balance of lease payable - reduction in lease payable)
Jan 1, 2021 $                              307,157
Dec 31, 2021 $          307,157 $        108,000 $                   18,429 $               89,571 $                              217,586
Dec 31, 2022 $          217,586 $        108,000 $                   13,055 $               94,945 $                              122,642
Dec 31, 2023 $          122,642 $        130,000 $                     7,358 $            122,642 $                                           0
Last payments includes = 108000 + 22000 = 130000
Lessee (Maywood )
Journal entries
Date General Journal Debit Credit
Jan 1, 2021 Right of use asset $        307,157
Lease liability $          307,157
(To record the lease)
Dec 31, 2021 Lease liability $          89,571
Interest expense $          18,429
Cash $          108,000
(To record lease payment)
Dec 31, 2021 Amortization Expense $        102,386
Right of use Asset $          102,386
(To record depreciation) (307157/3)